OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 5? a. x = 2 b. x = 4 c. x = 32 d. x = 128

OpenStudy (apoorvk):

\[\large log_ma - log_mb = log_m(a/b)\] Use this to club together the two terms in LHS, and then take anti-log both sides.

OpenStudy (anonymous):

ok so its log 2 x/4 = 5 now?

OpenStudy (apoorvk):

Yes!

OpenStudy (anonymous):

so what else?

OpenStudy (anonymous):

how do u get the answer?

OpenStudy (callisto):

\[log_2x-log_24 = 5\]\[log_2\frac{x}{4} = 5\]\[\frac{x}{4} = 2^5\]\[\frac{x}{4} = 32\]Can you solve x?

OpenStudy (anonymous):

why 2^5

OpenStudy (apoorvk):

because: